if cotA+1/cotA=2 then find the value of cot^2A+1/cot^2A

Grade:12

2 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Please find the answer to your question below
givencotA+1/cotA=2
let x=cotA
then x+1/x=2
squaring on both sides we get
x2+1/x2+2=4
so x2+1/x2=2
substituting x=cotA in above expression we get
cot2A+1/cot2A=2
rahul
9 years ago
the answer shall be 2